using Avagadro's law,i have to determine the cost in dollar if a single aluminum atom in a roll or aluminum foil. roll cost $3.29 for 6.96 sq meters (m2).
1
Expert's answer
2013-02-28T10:35:04-0500
Aluminium foil is alluminium prepared in thin metal leaves, with a thickness less than 0.2 millimetres =0.0002 meters.

The density of aluminum is 2700 kg/m3

Volume of foil is 6.96 m^2 * 0.0002 m =0,00139 m^3

Mass of it is 2700 kg/m3 *0,00139 m^3 =3,758

Amount of Al is 3,758/26.98 =0,1393 mol

Using Avogadro's law:

0,1393 N
Al-------------->Al
& 6.02 *10^23


N=0.84*10^23 atoms

Cost in dollar of a single aluminum atom is 3.29$/0.84*10^23 atoms = 3.91*10^-23